### 21 Jan 2024 v1

# Rings whose invertible elements are weakly nil-clean

Abstract.
We study those rings in which all invertible elements are weakly nil-clean calling them UWNC rings. This somewhat extends results due to Karimi-Mansoub et al. in Contemp. Math. (2018), where rings in which all invertible elements are nil-clean were considered abbreviating them as UNC rings. Specifically, our main achievements are that the triangular matrix ring ${\rm T}_{n}(R)$ over a ring $R$ is UWNC precisely when $R$ is UNC. Besides, the notions UWNC and UNC do coincide when $2$ inverts in $R$. We also describe UWNC $2$-primal rings $R$ by proving that $R$ is a ring with nil $J(R)$ such that $U(R)=\pm 1+{\rm Nil}(R)$. In particular, the polynomial ring $R[x]$ over some arbitrary variable $x$ is UWNC exactly when $R$ is UWNC. Some other relevant assertions are proved in the present direction as well.
Key words and phrases:
nil-clean element (ring), weakly nil-clean element (ring), UU-rings, weakly UU-rings
2010 Mathematics Subject Classification:
16S34, 16U60
1. Introduction and Major Concepts
Everywhere in the current paper, let $R$ be an associative but not necessarily commutative ring having identity element, usually stated as $1$. Standardly, for such a ring $R$, the letters $U(R)$, $\rm{Nil}(R)$ and ${\rm Id}(R)$ are designed for the set of invertible elements (also termed as the unit group of $R$), the set of nilpotent elements and the set of idempotent elements in $R$, respectively. Likewise, $J(R)$ denotes the Jacobson radical of $R$, and ${\rm Z}(R)$ denotes the center of $R$. The ring of $n\times n$ matrices over $R$ and the ring of $n\times n$ upper triangular matrices over $R$ are denoted by ${\rm M}_{n}(R)$ and ${\rm T}_{n}(R)$, respectively. Standardly, a ring is said to be abelian if each of its idempotents is central, that is, ${\rm Id}(R)\subseteq{\rm Z}(R)$.
For all other unexplained explicitly notions and notations, we refer to the classical source [15] or to the cited in the bibliography research sources. However, for completeness of the exposition and for the reader’s convenience, we recall the following basic notions.
Definition 1.1 ([13]).
Let $R$ be a ring. An element $r\in R$ is said to be nil-clean if there is an idempotent $e\in R$ and a nilpotent $b\in R$ such that $r=e+b$. Such an element $r$ is further called strongly nil-clean if the existing idempotent and nilpotent can be chosen such that $be=eb$. A ring is called nil-clean (respectively, strongly nil-clean) if each of its elements is nil-clean (respectively, strongly nil-clean).
Definition 1.2 ([12],[3],[6]).
A ring $R$ is said to be weakly nil-clean provided that, for any $a\in R$, there exists an idempotent $e\in R$ such that $a-e$ or $a+e$ is nilpotent. A ring $R$ is said to be strongly weakly nil-clean provided that, for any $a\in R$, $a$ or $-a$ is strongly nil-clean.
Definition 1.3 ([4],[11]).
A ring is called UU if all of its units are unipotent,that is, $U(R)\subseteq 1+{\rm Nil}(R)$ (and so $1+{\rm Nil}R)=U(R)$).
Definition 1.4 ([9]).
A ring $R$ is called weakly UU and abbreviated as $WUU$ if $U(R)={\rm Nil}(R)\pm 1$. This is equivalent to the condition that every unit can be presented as either $n+1$ or $n-1$, where $n\in{\rm Nil}(R)$.
Definition 1.5 ([14]).
A ring $R$ is called UNC if every of its units is nil-clean.
Our key working instrument is the following one.
Definition 1.6.
A ring $R$ is called UWNC if every of its units is weakly nil-clean.
In [14], the authors investigated UNC rings, i.e., those rings whose units are nil-clean. Our plan is to expand this substantially by developing an useful for this purpose machinery. In fact, we use some non-standard techniques from ring and matrix theories as well as also compare and distinguish the established results with these from [9].
It is worth noticing that some closely related background material can also be found in the sources [10] and [16].
The next constructions that show some proper ring classes inclusions are worthwhile.
Example 1.7.
  1. (i)
    Any nil-clean ring is weakly nil-clean, but the converse is not true in general. For instance, $\mathbb{Z}_{3}$ is weakly nil-clean but is not nil-clean.
  2. (ii)
    Any UU ring is WUU, but the converse is not true in general. For instance, $\mathbb{Z}$ is a WUU ring but is not UU.
  3. (iii)
    Any UU ring and nil-clean ring are UNC, but the converse is not true. For instance, the direct sum $\mathbb{Z}_{2}[t]\oplus{\rm M}_{2}(\mathbb{Z}_{2})$ is a UNC ring which is neither UU nor nil-clean.
  4. (iv)
    Any WUU ring and weakly nil-clean ring are UWNC, but the converse is not true in general. For instance, the direct sum $\mathbb{Z}\oplus M_{2}(\mathbb{Z}_{2})$ is a UWNC ring which is neither WUU nor weakly nil-clean.
  5. (v)
    Any UNC ring is UWNC, but the converse is not true in general. For instance, all of the rings $\mathbb{Z}_{3}$, $\mathbb{Z}_{6}$, $\mathbb{Z}_{3}\oplus{\rm M}_{2}(\mathbb{Z}_{2})$ are UWNC but are not UNC.
nil-clean weakly nil-clean UNC UWNC UU UNC WUU UWNC
Our further programm is the following: In the next second section, we obtain some main properties of the newly defined class of UWNC rings – the main results here are Proposition 2.12 as well as Theorems theo2.7, 2.19, 2.21 and 2.29, respectively. In the subsequent third section, we explore UWNC group rings – the main results here are Propositions 3.3 and 3.5, respectively.
2. Basic Properties of UWNC Rings
We start here with the following obvious technicality.
Proposition 2.1.
A unit $u$ of a ring $R$ is strongly weakly nil-clean if, and only if, $u\in\pm 1+{\rm Nil}(R)$. In particular, $R$ is a WUU ring if, and only if, every unit of $R$ is strongly weakly nil-clean.
We continue our work with the next two technical claims as follows.
Proposition 2.2.
Let $R$ be a UWNC ring and $S$ a UNC ring. Then, $R\times S$ is a UWNC ring.
Proof.
Choose $(u,v)\in U(R\times S)=U(R)\times U(S)$. Thus, there exist an idempotent $e\in R$ and a nilpotent $n\in R$ such that $u=e+n$ or $u=-e+n$. We now differ two cases:
Case I. Write $u=e+n$. Then, we have an idempotent $f\in S$ and a nilpotent $n^{\prime}\in S$ such that $v=f+n^{\prime}$. Thus, $(u,v)=(e,f)+(n,n^{\prime})$, where $(e,f)\in{\rm Id}(R\times S)$ and $(n,n^{\prime})\in{\rm Nil}(R\times S)$.
Case II. Write $u=-e+n$. Then, we have an idempotent $f\in S$ and a nilpotent $n^{\prime}\in S$ such that $-v=f+n^{\prime}$. Thus, $(u,v)=-(e,f)+(n,-n^{\prime})$, where $(e,f)\in{\rm Id}(R\times S)$ and $(n,-n^{\prime})\in{\rm Nil}(R\times S)$.
Therefore, $(u,v)$ is either the sum or difference of a nilpotent and an idempotent in $R\times S$, whence we get the desired result. ∎
Proposition 2.3.
Let $\{R_{i}\}$ be a family of rings. Then, the direct product $R=\prod R_{i}$ of rings $R_{i}$ is UWNC if, and only if, each $R_{i}$ is UWNC and at most one of them is not UNC.
Proof.
($\Rightarrow$). Obviously, each $R_{i}$ is UWNC. Suppose now $R_{i_{1}}$ and $R_{i_{2}}$ $(i_{1}\neq i_{2})$ are not UNC. Then, there exist some $u_{i_{j}}\in U(R_{i_{j}})\quad(j=1,2)$ such that $u_{i_{1}}\in U(R_{i_{1}})$ and $-u_{i_{2}}\in U(R_{i_{2}})$ are both not nil-clean. Choosing $u=(u_{i})$, where $u_{i}=0$ whenever $i\neq i_{j}\quad(j=1,2)$, we infer that $u$ and $-u$ are it not the sum of an idempotent and a nilpotent, as required to get a contradiction. Consequently, each $R_{i}$ is a UWNC ring as at most one of them is not UNC.
($\Leftarrow$). Assume that $R_{i_{0}}$ is a UWNC ring and all the others $R_{i}$ are UNC. So, a simple check gives that $\prod_{i\neq i_{0}}R_{i}$ is UNC. According to Proposition 2.2, we conclude that $R$ is a UWNC ring. ∎
However, the property of being UWNC is not closed under taking (internal, external) direct sum as the next construction illustrate.
Example 2.4.
A ring $\mathbb{Z}_{3}$ is a UWNC ring, but $\mathbb{Z}_{3}\times\mathbb{Z}_{3}$ is not UWNC.
Three further helpful affirmations are the following.
Corollary 2.5.
Let $L=\prod_{i\in I}R_{i}$ be the direct product of rings $R_{i}\cong R$ and $|I|\geq 2$. Then, $L$ is a UWNC ring if, and only if, $L$ is a UNC ring if, and only if, $R$ is a UNC ring.
Corollary 2.6.
For any $n\geq 2$, the ring $R^{n}$ is UWNC if, and only if, $R^{n}$ is UNC if, and only if, $R$ is UNC.
Proposition 2.7.
Let $R$ be a UWNC ring. If $T$ is a factor-ring of $R$ such that all units of $T$ lift to units of $R$, then $T$ is a UWNC ring.
Proof.
Suppose that $f:R\rightarrow T$ is a surjective ring homomorphism. Let $v\in U(T)$. Then, there exists $u\in U(R)$ such that $v=f(u)$ and $u=\pm e+n$, where $e\in{\rm Id}(R)$ and $n\in{\rm Nil}(R)$. Therefore, we have $v=\pm f(e)+f(n)$, where $f(e)\in{\rm Id}(T)$ and $f(n)\in{Nil}(T)$, as needed. ∎
We now offer the validity of the following statement.
Theorem 2.8.
Let $R$ be a ring and $I$ a nil-ideal of $R$.
  1. (i)
    $R$ is a UWNC ring if, and only if, $J(R)$ is nil and $\dfrac{R}{J(R)}$ is a UWNC ring.
  2. (ii)
    $R$ is a UWNC ring if, and only if, $\dfrac{R}{I}$ is a UWNC ring.
Proof.
  1. (i)
    Let $R$ be a UWNC ring and suppose $x\in J(R)$ and $x\notin{\rm Nil}(R)$. Since $1+x\in U(R)$, it must be that $1+x=-e+n$, where $n\in{\rm Nil}(R)$ and $e\in{\rm Id}(R)$, because if $1+x=e+n$, we will have $x\in{\rm Nil}(R)$ that is a contradiction.
    So, $2+x\in{\rm Nil}(R)$. Similarly, since $1+x^{2}\in U(R)$, we deduce that $2+x^{2}\in{\rm Nil}(R)$. Hence,
    $(2+x^{2})-(2+x)=x^{2}-x=-x(1-x)\in{\rm Nil}(R).$
    But $1-x\in U(R)$ whence $x\in{\rm Nil}(R)$, a contradiction. Thus, $J(R)$ is nil.
    Now, letting $\bar{u}\in U\left(\bar{R}=\dfrac{R}{J(R)}\right)$, we obtain $u\in U(R)$, because units lift module $J(R))$. Therefore, $u=\pm e+n$, where $e\in{\rm Id}(R)$ and $n\in{\rm Nil}(R)$. So, we have $\bar{u}=\pm\bar{e}+\bar{n}$, where $\bar{e}\in{\rm Id}(\bar{R})$ and $\bar{n}\in{\rm Nil}(\bar{R})$. Thus, $\dfrac{R}{J(R)}$ is a UWNC ring, as promised.
    Conversely, let $u\in U(R)$. Then, $\bar{u}\in U(\bar{R})$ and write $\bar{u}=\pm\bar{e}+\bar{n}$, where $\bar{e}\in{\rm Id}(\bar{R})$ and $\bar{n}\in{\rm Nil}(\bar{R})$. As $J(R)$ is nil, idempotents of $\dfrac{R}{J(R)}$ can be lifted to idempotents of $R$. So, we can assume that $e^{2}=e\in R$. Moreover, one inspects that $n\in R$ is nilpotent. Thus, for some $j\in J(R)$,
    $u=\pm e+n+j=\pm e+(n+j)$
    is weakly nil-clean, because $n+j\in{\rm Nil}(R)$, as expected.
  2. (ii)
    The proof is similar to (i), so we omit the details.
As an immediate consequence, we yield:
Corollary 2.9.
Let $R$ be a ring and $M$ a bi-module over $R$. Then, the following hold:
  1. (i)
    The trivial extension ${\rm T}(R,M)$ is a UWNC ring if, and only if, $R$ is a UWNC ring.
  2. (ii)
    For $n\geq 2$, the quotient-ring $\dfrac{R[x]}{\langle x^{n}\rangle}$ is a UWNC ring if, and only if, $R$ is a UWNC ring.
  3. (iii)
    For $n\geq 2,thequotient-ring\dfrac{R[[x]]}{\langle x^{n}\rangle}$ is a UWNC ring if, and only if, $R$ is a UWNC ring.
Proof.
  1. (i)
    Set $A={\rm T}(R,M)$ and consider $I:={\rm T}(0,M)$. It is not too hard to verify that $I$ is a nil-ideal of $A$ such that $\dfrac{A}{I}\cong R$. So, the result follows directly from Theorem 2.8.
  2. (ii)
    Put $A=\dfrac{R[x]}{\langle x^{n}\rangle}$. Considering $I:=\dfrac{\langle x\rangle}{\langle x^{n}\rangle}$, we obtain that $I$ is a nil-ideal of $A$ such that $\dfrac{A}{I}\cong R$. So, the result follows automatically Theorem 2.8.
  3. (iii)
    Knowing that the isomorphism $\dfrac{R[x]}{\langle x^{n}\rangle}\cong\dfrac{R[[x]]}{\langle x^{n}\rangle}$ is true, point (3) follows automatically from (2).
Another two consequences of interest are the following ones:
Corollary 2.10.
Let $R$, $S$ be rings and let $M$ be an $(R,S)$-bi-module. If $T=\begin{pmatrix}R&M\\ 0&S\end{pmatrix}$ is a UWNC ring, then both $R,S$ are UWNC rings. The converse holds provided one of the rings $R$ or $S$ is UNC and the other is UWNC.
Proof.
Given $I:=\left(\begin{array}[]{ll}0&M\\ 0&0\end{array}\right)$. A routine inspection shows that this is a nil-ideal in $T$ with $\dfrac{T}{I}\cong R\times S$. Therefore, the result follows applying Proposition 2.3 and Theorem 2.8. ∎
Corollary 2.11.
Let $R$ be a ring. Then, the following are equivalent:
  1. (i)
    $R$ is a UWNC ring.
  2. (ii)
    ${\rm T}_{n}(R,\alpha)$ is a UWNC ring.
Proof.
Choose
$I:=\left\{\left.\begin{pmatrix}0&a_{12}&\ldots&a_{1n}\\ 0&0&\ldots&a_{2n}\\ \vdots&\vdots&\ddots&\vdots\\ 0&0&\ldots&0\end{pmatrix}\right|a_{ij}\in R\quad(i\leq j)\right\}.$
Then, one easily verifies that $I^{n}=0$ and $\dfrac{{\rm T}_{n}(R,\alpha)}{I}\cong R$. Consequently, Theorem 2.8 applies to get the wanted result. ∎
The next assertion is pivotal.
Proposition 2.12.
Let $R$ be a ring. Then, the following are equivalent:
  1. (i)
    $R$ is a UNC ring.
  2. (ii)
    ${\rm T}_{n}(R)$ is a UNC ring for all $n\in\mathbb{N}$.
  3. (iii)
    ${\rm T}_{n}(R)$ is a UNC ring for some $n\in\mathbb{N}$.
  4. (iv)
    ${\rm T}_{n}(R)$ is a UWNC ring for some $n\geq 2$.
Proof.
(1) $\Rightarrow$ (2). This follows employing [14, Corollary 2.6].
(2) $\Rightarrow$ (3) $\Rightarrow$ (4). These two implications are trivial, so we remove the details.
(4) $\Rightarrow$ (1). Method 1: Let $u\in U(R)$ and choose
$A=\begin{pmatrix}u&&&\ast\\ &-u_{1}&&\\ &&\ddots&\\ 0&&&1\end{pmatrix}\in U({\rm T}_{n}(R)).$
By hypothesis, we can find an idempotent $\begin{pmatrix}e_{1}&&&\ast\\ &e_{2}&&\\ &&\ddots&\\ 0&&&e_{n}\end{pmatrix}$ and a nilpotent
$\begin{pmatrix}w_{1}&&&\ast\\ &w_{2}&&\\ &&\ddots&\\ 0&&&w_{n}\end{pmatrix}$ such that
$A=\pm\begin{pmatrix}e_{1}&&&\ast\\ &e_{2}&&\\ &&\ddots&\\ 0&&&e_{n}\end{pmatrix}+\begin{pmatrix}w_{1}&&&\ast\\ &w_{2}&&\\ &&\ddots&\\ 0&&&w_{n}\end{pmatrix}.$
It now follows that $u=e_{1}+w_{1}$ or $u=e_{2}-w_{2}$. Clearly, $e_{1}$, $e_{2}$ are idempotents and $w_{1}$, $w_{2}$ are nilpotents in $R$, thus proving point (1).
Method 2: Setting $I:=\{(a_{ij})\in T_{n}(R)|a_{ii}=0\}$, we obtain that it is a nil-ideal in ${\rm T}_{n}(R)$ with $\dfrac{{\rm T}_{n}(R)}{I}\cong R^{n}$. Therefore, Theorem 2.8 and Corollary 2.6 are applicable to get the pursued result. ∎
We know that $\mathbb{Z}\oplus{\rm M}_{2}(\mathbb{Z}_{2})$ is a UWNC ring that is neither WUU nor weakly nil-clean. The following example concretely demonstrates an indecomposable UWNC ring that is neither WUU nor weakly nil-clean.
Example 2.13.
Let $R={\rm M}_{2}(\mathbb{Z}_{2})$, $S=\mathbb{Z}_{3}[x]$ and $M=\{0\}$. Then, $M$ is obviously an $(R,S)$-bi-module, and the formal triangular matrix ring $T:=\begin{pmatrix}R&M\\ 0&S\end{pmatrix}$ is a UWNC ring using Corollary 2.10. Furthermore, one can verify that the central idempotents of $T$ are trivial. So, $T$ is necessarily indecomposable. But since $R$ is not a WUU ring and $S$ is not a weakly nil-clean ring, one plainly follows that $T$ is neither a WUU ring nor a weakly nil-clean ring, as claimed.
It was proved in [9, Proposition 2.25] that any unital subring of a WUU ring is again a WUU ring. But, curiously, a subring of a UWNC ring may not be a UWNC ring as the next example shows.
Example 2.14.
Let $T={\rm M}_{2}(\mathbb{Z}_{2})$ and $u=\begin{pmatrix}0&1\\ 1&1\end{pmatrix}\in T$. Then, one sees that $u^{3}=1$. Now, let $R$ be the unital subring of $T$ generated by $u$. Therefore, one calculates that
$R=\{a1+bu+cu^{2}|a,b,c\in\mathbb{Z}\}=\{0,1,u,u^{2},1+u,1+u^{2},1+u+u^{2},u+u^{2}\}.$
But, we have that $1+u=u^{2}$, $1+u^{2}=u$, $01+u+u^{2}=$ 0 and $u+u^{2}=1$, so we deduce $R=\{0,1,u,u^{2}\}$. It is now easy to see that ${\rm Nil}(R)=\{0\}$, so that $R$ is reduced. As $u^{2}\neq u$, $u$ is manifestly not weakly nil-clean utilizing [3, Theorem 20], and so $T$ is a UWNC ring, as asserted.
Proposition 2.15.
The following two statements are valid:
  1. (i)
    If $R$ is a weakly nil-clean ring, then ${\rm Z}(R)$ is strongly weakly nil-clean.
  2. (ii)
    If $R$ is a UWNC ring, then ${\rm Z}(R)$ is a WUU ring.
Proof.
  1. (i)
    Let $a\in{\rm Z}(R)$. Then, $a\in R$ is weakly nil-clean and central, so $a$ is strongly weakly nil-clean in $R$. Thus, $a\pm a^{2}\in{\rm Nil}(R)$ by [5, Theorem 2.1]. But, $a\pm a^{2}\in{\rm Z}(R)$ so that
    $a\pm a^{2}\in{\rm Nil}(R)\cap{\rm Z}(R)\subseteq{\rm Nil}({\rm Z}(R)).$
    Hence, ${\rm Z}(R)$ is strongly weakly nil-clean, as required.
  2. (ii)
    The proof is analogous to (1).
Proposition 2.16.
For any ring $R$, the power series ring $R[[x]]$ is not UWNC.
Proof.
Note the principal fact that the Jacobson radical of $R[[x]]$ is not nil (see, e.g., [15]). Thus, in view of Theorem 2.8, $R[[x]]$ is really not a UWNC ring, as expected. ∎
Lemma 2.17.
Let $R$ be a ring. Then, the following two points are equivalent:
  1. (i)
    $R$ is a UNC ring.
  2. (ii)
    $R$ is a UWNC ring and $2\in J(R)$.
Proof.
(1) $\Longrightarrow$ (2). Evidently, $R$ is a UWNC ring. Also, we have $2\in J(R)$ in virtue of [14, Lemma 2.4].
(2) $\Rightarrow$ (1). Notice that $\dfrac{R}{J(R)}$ is of characteristic $2$, because $2\in J(R)$), and so $a=-a$ for every $a\in\dfrac{R}{J(R)}$. That is why, $\dfrac{R}{J(R)}$ is a UNC ring, and thus we can apply [14, Theorem 2.5] since $J(R)$ is nil in view of Theorem 2.8. ∎
Recall that an element $r$ in a ring $R$ is said to be an unipotent if $r-1$ is a nilpotent. The following technical claim is elementary, but rather applicable in the sequel.
Lemma 2.18.
Let $R$ be a ring and let $r\in R$ be the sum of an idempotent and a nilpotent. If $r^{2}=1$, then $r$ is unipotent.
Proof.
Write $r=e+n$ with $e\in{\rm Id}(R)$ and $n\in{\rm Nil}(R)$. Set $f:=1-e$ and $x:=n(n+1)\in{\rm Nil}(R)$. Taking into account the equality $fn=f(r-e)=fr$, we compute that
$fx=fn(n+1)=fr(r-e+1)=fr(r+f)=fr^{2}+frf=f+frf,$
and, similarly, that $xf=f+frf$. Hence, $fx=xf$, so that $x$ is a nilpotent which commutes with $f$, $e$, $n$ and $r$, respectively. Accordingly,
$f=fr^{2}=fr\cdot r=fnr=fx(1+n)^{-1}r=f(1+n)^{-1}r\cdot x$
is a nilpotent and hence $f=0$, as desired. ∎
Our next main result, which sounds quite surprisingly, is the following.
Theorem 2.19.
Let $R$ be a ring and $2\in U(R)$. Then, the following two items are equivalent:
  1. (i)
    $R$ is a UWNC ring.
  2. (ii)
    $R$ is a WUU ring.
Proof.
(2) $\Rightarrow$ (1). This is pretty obvious, so we leave the argumentation.
(1) $\Rightarrow$ (2). First, we show that $R$ is an abelian ring. To this goal, let $e^{2}=e\in R$, and let $a=1-2e$. Then, it is obviously true that $a^{2}=1$. Since $R$ is UWNC, either $a$ or $-a$ is nil-clean. By virtue of Lemma 2.18, one has that $a\in 1+{\rm Nil}(R)$ or $a\in-1+{\rm Nil}(R)$. If $a\in 1+{\rm Nil}(R)$, then $2e\in{\rm Nil}(R)$, and so $e\in{\rm Nil}(R)$. This implies that $e=0$. If, however, $a\in-1+{\rm Nil}(R)$, then $2(1-e)\in{\rm Nil}(R)$, whence $1-e\in{\rm Nil}(R)$. This forces that $e=1$. Therefore, $R$ has only trivial idempotents. Thus, $R$ is abelian, as asserted.
Now, let $u\in U(R)$, so $u=\pm e+n$, where $e\in{\rm Id}(R)$ and $n\in{\rm Nil}(R)$. If $u=e+n$, so $e=u-n\in U(R)$, then $e=1$. If, however, $u=-e+n$, so $e=u-n\in U(R)$, then $e=1$. Therefore, $u\in\pm 1+{\rm Nil}(R)$. Finally, one concludes that $R$ is WUU, as formulated. ∎
As an immediate consequence, we derive:
Corollary 2.20.
Let $R$ be an abelian ring. Then, the following are equivalent:
  1. (i)
    $R$ is a UWNC ring.
  2. (ii)
    $R$ is a WUU ring.
We now arrive at our third chief result.
Theorem 2.21.
Let $R$ be a 2-primal ring and $\alpha$ an endomorphism of $R$ such that $R$ is $\alpha$-compatible. The following issues are equivalent:
  1. (i)
    $R[x,\alpha]$ is a UWNC ring.
  2. (ii)
    $R[x,\alpha]$ is a WUU ring.
  3. (iii)
    $R$ is a WUU ring.
  4. (iv)
    $R$ is a UWNC ring.
  5. (v)
    $J(R)={\rm Nil}(R)$ and $U(R)=1\pm J(R)$.
Proof.
(ii) $\Rightarrow$ (i) and (iii) $\Rightarrow$ (iv). Straightforward.
(i) $\Rightarrow$ (iv). As $\dfrac{R[x,\alpha]}{\langle x\rangle}\cong R$ and all units of $\dfrac{R[x,\alpha]}{\langle x\rangle}$ are lifted to units of $R[x,\alpha]$, the implication easily holds.
(ii) $\Longrightarrow$ (iii). We argue as in the proof of (1) $\Longrightarrow$ (4).
(iv) $\Rightarrow$ (v). As $R$ is 2-primal, we have ${\rm Nil}(R)\subseteq J(R)$, so that $J(R)={\rm Nil}(R)$ bearing in mind Theorem 2.8. Let $a\in U(R)$, so by hypothesis we have $u=\pm e+n$, where $e\in{\rm Id}(R)$ and $n\in{\rm Nil}(R)=J(R)$. If $u=e+n$, so $e=u-n\in U(R)$, and thus $e=1$. If, however, $u=-e+n$, so $e=n-u\in U(R)$, and thus $e=1$. Therefore, we receive $u\in\pm 1+J(R)$, and hence $U(R)=\pm 1+J(R)$, as required.
(v) $\Rightarrow$ (ii). As $R$ is a $2$-primal ring, with the aid of (5) we have $J(R)={\rm Nil}(R)=\mathrm{Nil}_{\ast}(R)$. Thus, the quotient-ring $\dfrac{R}{J(R)}$ is a reduced ring. Moreover, it is easy to see that $\alpha({\rm Nil}(R))\subseteq{\rm Nil}(R)$, so $\alpha(J(R))\subseteq J(R)$ and $\bar{\alpha}:\dfrac{R}{J(R)}\rightarrow\dfrac{R}{J(R)}$, defined by $\bar{\alpha}(\bar{a})=\overline{\alpha(a)}$, is an endomorphism of $\dfrac{R}{J(R)}$.
We next show that $\dfrac{R}{J(R)}$ is $\bar{\alpha}$-compatible. That is, we must show that, for any $a+J(R),b+J(R)\in\dfrac{R}{J(R)}$, the equivalence
$(a+J(R))(b+J(R))=J(R)\Leftrightarrow(a+j(R))\bar{\alpha}(b+j(R))=J(R)$
holds. Equivalently, we have to show that, for any $a,b\in R$, the equivalence $ab\in{\rm Nil}(R)\Leftrightarrow a\alpha(b)\in{\rm Nil}(R)$ is true. But this equivalence has been established in the proof of Claims 1 and 2 in [1, Theorem 3.6]. As $\dfrac{R}{J(R)}$ is a reduced factor-ring and also is $\bar{\alpha}$-compatible, with [7, Corollary 2.12] at hand we have
$U\left(\dfrac{R}{J(R)}[x,\bar{\alpha}]\right)=U\left(\dfrac{R}{J(R)}\right),$
which is equal to {±$\{\pm\bar{1}\}$ by assumption. So,
$\dfrac{R}{J(R)}[x,\bar{\alpha}]\cong\dfrac{R[x,\alpha]}{J(R)[x,\alpha]}$
is a WUU ring. Also, [7, Lemma 2.2] tells us that
$\mathrm{Nil}_{*}(R[x,\alpha])=\mathrm{Nil}_{*}(R)[x,\alpha].$
Therefore,
$J(R)[x,\alpha]=\mathrm{Nil}_{*}(R[x,\alpha]),$
which is manifestly nil. Hence, [9, Proposition 2.5 and 2.6] ensures that $R[x,\alpha]$ is a WUU ring, as asked for. ∎
As a direct consequence, we deduce:
Corollary 2.22.
Let $R$ be a 2-primal ring. Then, the following are equivalent:
  1. (i)
    $R$ is a UWNC ring.
  2. (ii)
    $R[x]$ is a UWNC ring.
  3. (iii)
    $J(R)={\rm Nil}(R)$ and $U(R)=\pm 1+J(R)$.
The following criterion is worthy of documentation.
Proposition 2.23.
Suppose $R$ is a commutative ring. Then, $R[x]$ is a UWNC ring if, and only if, $R$ is a UWNC ring.
Proof.
Method 1: For the necessity, let $R[x]$ be a UWNC ring. We know that $\dfrac{R[x]}{\langle x\rangle}\cong R$. It, therefore, suffices to show that $\dfrac{R[x]}{\langle x\rangle}$ is a UWNC ring. To this aim, choosing $u+\langle x\rangle\in U\left(\dfrac{R[x]}{\langle x\rangle}\right)$, we derive $u\in U(R[x])$, because all units of $\dfrac{R[x]}{\langle x\rangle}$ are lifted to units of $R[x]$. Thus, we write $u=\pm e+n$, where $e\in{\rm Id}(R[x])$ and $n\in{\rm Nil}(R[x])$. So,
$u+\langle x\rangle=\pm(e+\langle x\rangle)+(n+\langle x\rangle),$
where $e+\langle x\rangle\in{\rm Id}\left(\dfrac{R[x]}{\langle x\rangle}\right)$ and $n+\langle x\rangle\in{\rm Nil}\left(\dfrac{R[x]}{\langle x\rangle}\right)$. Consequently, $u+\langle x\rangle$ is a weakly nil-clean element and $\dfrac{R[x]}{\langle x\rangle}$ is a UWNC ring, as desired.
For the sufficiency, write
$f=a_{0}+a_{1}x+\ldots+a_{n}x^{n}\in U(R[x]),$
so $a_{0}\in U(R)$ and $a_{1},\ldots,a_{n}\in{\rm Nil}(R)$. By hypothesis, we have $a_{0}=\pm e+n$, where $e\in{\rm Id}(R)$ and $n\in{\rm Nil}(R)$. This allows us to infer that
$f=(\pm e+n)+a_{1}x+\cdots+a_{n}x^{n}=\pm e+(n+a_{1}x+\cdots+a_{n}x^{n}),$
where $e\in{\rm Id}(R[x])$ and $n+a_{1}x+\cdots+a_{n}x^{n}\in{\rm Nil}(R[x])$. Finally, $f$ is a weakly nil-clean element, thus establishing the result.
Method 2: It is well know that every commutative ring is a $2$-primal ring, and hence the result can be deduced from Corollary 2.22. ∎
Incidentally, we are able to prove the following curious statement.
Proposition 2.24.
Let $R$ be a ring, and $m,n\geq 1$. If the matrix rings ${\rm M}_{n}(R)$ and ${\rm M}_{m}(R)$ are both UWNC, then so is the triangular matrix ring ${\rm T}_{n+m}(R)$.
Proof.
Let $V\in U({\rm T}_{n+m}(R))$ be the $(n+m)\times(n+m)$ triangular matrix which we will write in the block decomposition form as follows $V=\begin{pmatrix}V_{11}&A_{12}\\ 0&V_{22}\end{pmatrix}$, where $V_{11}\in U({\rm M}_{n}(R))$, $V_{22}\in U({\rm M}_{m}(R))$ and $A_{12}$ is appropriately sized rectangular matrices. By hypothesis, there exist idempotent matrices $E_{1}$, $E_{2}$ and nilpotent matrices $N_{1}$, $N_{2}$ in ${\rm M}_{n}(R)$ and ${\rm M}_{m}(R)$ such that $V_{11}=\pm E_{1}+N_{1}$ and $V_{22}=\pm E_{2}+N_{2}$. Thus, we obtain the decomposition
$\begin{pmatrix}V_{11}&A_{12}\\ 0&V_{22}\end{pmatrix}$ =(±$=\begin{pmatrix}\pm E_{1}+N_{1}&A_{12}\\ 0&\pm E_{2}+N_{2}\end{pmatrix}$
=±$=\pm\begin{pmatrix}E_{1}&0\\ 0&E_{2}\end{pmatrix}+\begin{pmatrix}N_{1}&A_{12}\\ 0&N_{2}\end{pmatrix}.$
Since $E_{1}$, $E_{2}$ are idempotents, then an easy verification guarantees that $\begin{pmatrix}E_{1}&0\\ 0&E_{2}\end{pmatrix}$ is an idempotent. It is also readily to see that $\begin{pmatrix}N_{1}&A_{12}\\ 0&N_{2}\end{pmatrix}$ is a nilpotent. Thus, the above decomposition is the desired weakly nil-clean decomposition. ∎
Proposition 2.25.
Let $R=\left(\begin{array}[]{ll}A&M\\ N&B\end{array}\right)$ be a Morita context ring such that $MN$ and $NM$ are nilpotent ideals of $A$ and $B$, respectively. If $R$ is a UWNC ring, then $A$ and $B$ are UWNC rings. The converse holds provided one of the $A$ or $B$ is UNC and the other is UWNC.
Proof.
We know that $J(R)=\left(\begin{array}[]{cc}J(A)&M_{0}\\ N_{0}&J(B)\end{array}\right)$, where $M_{0}=\{x\in M\mid xN\subseteq J(A)\}$ and $N_{0}=\{y\in N\mid yM\subseteq J(B)\}$ in accordance with [2, Lemma 4.4]. As $MN$ and $NM$ are nilpotent ideals of $A$ and $B$, respectively, we therefore have that $\dfrac{R}{J(R)}\cong\dfrac{A}{J(A)}\times\dfrac{B}{J(B)}$. Since $R$ is a UWNC ring, then the factor $\dfrac{R}{J(R)}$ is a UWNC ring and $J(R)$ is nil consulting with Theorem 2.8, so it follows that both $\dfrac{A}{J(A)}$ and $\dfrac{B}{J(B)}$ are UWNC. As $J(R)$ is nil, $J(A)$ and $J(B)$ are nil too. Thus, $A$ and $B$ are UWNC as well.
Oppositely, assuming that $A$ is a UNC ring and $B$ is a UWNC ring, we conclude that $\dfrac{R}{J(R)}$ is a UWNC ring by a combination of [14, Theorem 2.5], Theorem 2.8 and Proposition 2.2. It then suffices to show that $J(R)$ is nil. To this target, suppose $r=\begin{pmatrix}a&m\\ n&b\end{pmatrix}\in J(R)$ such that $a^{n}=b^{n}=0$ for some $n\in\mathbb{N}$. Then, a simple check assures that $r^{n}\in\begin{pmatrix}MN&M\\ N&NM\end{pmatrix}$ and, in particular, that
$\begin{pmatrix}MN&M\\ N&NM\end{pmatrix}=\begin{pmatrix}(MN)^{k}&(MN)^{k}M\\ (NM)^{k}N&(NM)^{k}\end{pmatrix}.$
Therefore, $J(R)$ is indeed nil, as desired. ∎
As three corollaries, we extract:
Corollary 2.26.
Let $R$ be a ring and $s\in{\rm Z}(R)\cap{\rm Nil}(R)$. If ${\rm K}_{s}(R)$ is a UWNC ring, then $R$ is a UWNC ring. The converse holds if $R$ is a UNC ring.
Corollary 2.27.
Let $R$ be a ring and $s\in{\rm Z}(R)\cap{\rm Nil}(R)$. If ${\rm M}_{n}(R;s)$ is a UWNC ring, then $R$ is a UWNC ring. The converse holds if $R$ is a UNC ring.
Corollary 2.28.
If the trivial Morita context $\begin{pmatrix}A&M\\ N&B\end{pmatrix}$ is a UWNC ring, then $A$, $B$ are UWNC rings. The converse holds if one of the rings $A$ or $B$ is UWNC and the other is UNC.
Proof.
It is apparent to see that the isomorphisms
$\begin{pmatrix}A&M\\ N&B\end{pmatrix}\cong{\rm T}(A\times B,M\oplus N)\cong\begin{pmatrix}A\times B&M\oplus N\\ 0&A\times B\end{pmatrix}$
are fulfilled. Then, the rest of the proof follows combining Corollary 2.9 and Proposition 2.3. ∎
We now intend to prove the following.
Theorem 2.29.
Let $R$ be a local ring. Then, the following are equivalent:
  1. (i)
    $R$ is a UWNC ring.
  2. (ii)
    $R$ is a weakly nil-clean ring.
Proof.
(1) $\Rightarrow$ (2). As $R$ is local, one finds that ${\rm Id}(R)=\{0,1\}$, and so $R$ is abelian. Therefore, $R$ is WUU owing to Corollary 2.20. Also, $R$ is clean, because every local ring is clean. Thus, $R$ is a clean WUU and we apply [9, Corollary 2.15] to find that $R$ is strongly weakly nil-clean and so it is weakly nil-clean.
(2) $\Rightarrow$ (1). It is clear. ∎
We say that $B$ is a unital subring of a ring $A$ if $\emptyset\neq B\subseteq A$ and, for any $x,y\in B$, the relations $x-y$, $xy\in B$ and $1_{A}\in B$ hold. Let $A$ be a ring and $B$ a unital subring of $A$, and denote by $R[A,B]$ the set $\{(a_{1},\ldots,a_{n},b,b,\ldots):a_{i}\in A,b\in B,1\leq i\leq n\}$. Then, $R[A,B]$ forms a ring under the usual component-wise addition and multiplication.
We, thereby, establish the following.
Proposition 2.30.
Let $A$ be a ring and a unital subring $B$ of $A$. Then, the following are equivalent:
  1. (i)
    $A$ and $B$ are UWNC.
  2. (ii)
    $R[A,B]$ is UWNC.
3. UWNC Group Rings
We begin here with the following simple but useful technicality.
Lemma 3.1.
Let $R$ and $S$ be rings and $i:R\rightarrow S$, $\varepsilon:S\rightarrow R$ be ring homomorphisms such that $\varepsilon i=id_{R}$.
  1. (i)
    $\varepsilon({\rm Nil}(S))={\rm Nil}(R)$, $\varepsilon(U(S))=U(R)$ and $\varepsilon({\rm Id}(S))={\rm Id}(R)$.
  2. (ii)
    If $S$ is a UWNC ring, then $R$ is a UWNC ring.
  3. (iii)
[…]
Proof.
  1. (i)
    Clearly, the inclusions $\varepsilon({\rm Nil}(S))\subseteq{\rm Nil}(R)$, $\varepsilon(U(S))\subseteq U(R)$ and $\varepsilon({\rm Id}(S))\subseteq{\rm Id}(R)$ are valid. On the other hand, we also have that ${\rm Nil}(R)=\varepsilon i({\rm Nil}(R))\subseteq\varepsilon({\rm Nil}(S))$, $U(R)=\varepsilon i(U(R))\subseteq\varepsilon(U(S))$ and ${\rm Id}(R)=\varepsilon i({\rm Id}(R))\subseteq\varepsilon({\rm Id}(S))$.
  2. (ii)
    Method 1: Let $S$ be a UWNC ring. Choose $u\in U(R)=\varepsilon(U(S))$, so $u=\varepsilon(v)$, where $v\in U(S)$. Thus, we can write $v=\pm e+q$, where $e\in{\rm Id}(S)$, $q\in{\rm Nil}(S)$. Therefore,
    $u=\varepsilon(v)=\varepsilon(\pm e+q)=\pm\varepsilon(e)+\varepsilon(q),$
    where $\varepsilon(e)\in{\rm Id}(R)$ and $\varepsilon(q)\in{\rm Nil}(R)$ exploiting (1).
    Method 2: Let $S$ be a UWNC ring. Hence, $U(S)=\pm{\rm Id}(S)+{\rm Nil}(S)$, whence by (1) one has that
    $U(R)=\varepsilon(U(S))=\varepsilon(\pm{\rm Id}(S)+{\rm Nil}(S))=\pm\varepsilon({\rm Id}(S))+\varepsilon({\rm Nil}(S))=\pm{\rm Id}(R)+{\rm Nil}(R),$
    as required.
  3. (iii)
    If $R$ is a UWNC ring, point (1) enables us that
    $U(S)=\varepsilon^{-1}(U(R))=\varepsilon^{-1}(\pm{\rm Id}(R)+{\rm Nil}(R))=\pm{\rm Id}(S)+{\rm Nil}(S)+\text{ker}\varepsilon=\pm{\rm Id}(S)+{\rm Nil}(S),$
    as required.
Remark 3.2.
It is a routine technical exercise to see that Lemma 3.1 (2) and (3) is true also for WUU rings.
Suppose that $G$ is an arbitrary group and $R$ is an arbitrary ring. As usual, $RG$ stands for the group ring of $G$ over $R$. The homomorphism $\varepsilon:RG\rightarrow R$, defined by $\varepsilon(\displaystyle\sum_{g\in G}a_{g}g)=\displaystyle\sum_{g\in G}a_{g}$, is called the augmentation map of $RG$ and its kernel, denoted by $\Delta(G)$, is called the augmentation ideal of $RG$.
Proposition 3.3.
Let $R$ be a ring and $G$ a group. If $RG$ is a UWNC ring, then $R$ is a UWNC ring. The converse holds if $\Delta(G)\subseteq{\rm Nil}(RG)$.
Proof.
Let us consider the inclusion $i:R\rightarrow RG$, given by $i(r)=\displaystyle\sum_{g\in G}a_{g}g$, where $a_{1G}=r$ and $a_{g}=0$ provided $g\neq 1_{G}$. It is easy to check that the map $i$ is a ring monomorphism and thus $R$ can also be viewed as a subring of $RG$. Furthermore, it is only enough to apply Lemma 3.1 (2) to get the claim. ∎
Corollary 3.4.
Let $R$ be a ring and $G$ a group. If $RG$ is a WUU ring, then $R$ is a WUU ring. The converse holds if $\Delta(G)\subseteq{\rm Nil}(RG)$.
Proof.
It follows at once by Proposition 3.3 and Lemma 3.1. ∎
Proposition 3.5.
Let $R$ be a UWNC ring with $p\in{\rm Nil}(R)$ and let $G$ be a locally finite $p$-group, where $p$ is a prime. Then, $RG$ is a UWNC ring.
Proof.
Referring to [8, Proposition 16], one verifies that $\Delta(G)$ is nil. Now, the assertion follows from the obvious isomorphism $\dfrac{RG}{\Delta(G)}\cong R$ and Theorem 2.8. ∎
Remark 3.6.
It is easily seen that Proposition 3.5 holds also for WUU rings.
Funding: The work of the first-named author, P.V. Danchev, is partially supported by the project Junta de Andalucía under Grant FQM 264, and by the BIDEB 2221 of TÜBÍTAK.